OpenStudy (anonymous):

A randomly mating population has the dominant allele “B” and the recessive allele “b” at the frequencies 0.7 and 0.3 respectively. A sudden cataclysmic event causes all the homozygous recessive individuals to die. However, the population simply carries on with life and gets busy mating. What was the proportion of homozygous recessive individuals before the cataclysm, and what is the proportion of homozygous recessive individuals born in the first generation after the cataclysm?

OpenStudy (blues):

Do you know the Castle-Hardy-Weinburg equation?

OpenStudy (blues):

You will need to use it.

OpenStudy (anonymous):

is it the p^2 +2pq + q^2=1 equation

OpenStudy (anonymous):

how would i start this question im a bit confused :S

OpenStudy (anonymous):

i know were trying to find q^2

OpenStudy (blues):

OK. You are given the frequencies p = 0.70 and q = 0.30 which you use in the CHW equation. The first thing you probably want to do at this point is find the frequencies of the different genotypes. That is: freq(pp) = 0.7^2 = 0.49 freq(2pq) = 2*0.7*0.3 = 0.42 freq(qq) = 0.3^2 = 0.09. Those are the frequencies before all the qq individuals die off. Next you want to calculate the frequencies of q and p after the die off.

OpenStudy (blues):

Do you have ideas about how to go about that?

OpenStudy (anonymous):

i wish i did lol

OpenStudy (blues):

OK, so after all the recessive homozygous individuals (qq) are dead, the only ones left are homozygous dominant (pp) and heterozygous (2pq) individuals. So your CHW equation, which gives the frequencies of each genotype in the population goes from this: p^2 + 2pq + q^2 = 1 to this: p^2 + 2pq = 1.

OpenStudy (blues):

The key is that removing all the qq individuals has changed the relative frequencies of the p and q alleles - there are more ps relative to fewer qs in the entire population.

OpenStudy (anonymous):

I hope you dont mind @blues. @jatt3092 asked for my help on the chat, so I'm just going to throw in a clue to help out. 'q' represents the allele frequency for the reccesive gene but it says that all the homo recessive individs die. But you can still have a recessive allele with a dominant individual. That would make them heterozygous, so go from there.

OpenStudy (blues):

Thanks Shay, the frequency equation after the die off (p^2 + 2pq = 1) plainly has the recessive allele (q) in it and the remark that the die off is merely a reduction of the recessive allele frequency relative to the dominant allele frequency has that clearly established. Out of curiosity, how would you go about solving the rest of the prob? : D

OpenStudy (anonymous):

im curious as well :D

OpenStudy (anonymous):

I'm actually not sure. q^2 should stay the same right? because the frequency of the population remains the same I think even if the allele is reduced. Or wait, can you multiply the q^2 by the total population to find the number of homo recessive individs? Then subtract that from the original total population which would be the new population because all the bb individs died. Then go from there, finding a new frequency for the p? I'm not too sure, but this might work.

OpenStudy (anonymous):

Or no, the new population would be the 2 pq and since the p^2, not q^2 remains the same, you can use math to find the rest I think

OpenStudy (blues):

That is, in a nutshell, what needs to be done next. Well done. :D You have the frequencies of the homozygous dominant and heterozygous individuals before the die off. Next you need to calculate their frequencies after the die off. It is easiest just to normalize by dividing by the frequency of remaining individuals. Frequency surviving = pp + 2pq. frequency surviving = 0.49 + 0.42 = 0.91. Find the post die of frequencies of homozygous dominant individuals and heterozygous individuals by dividing those frequencies by 0.91. So after the die off the frequencies of homozygous dominant and heterozygous individuals is this: pp = 0.49 / 0.91 = 0.5384 2pq = 0.42 / 0.91 = 0.46154. What happened to the person who we're supposedly tutoring?

OpenStudy (anonymous):

Hahaha... They got scared away I think. lol. I dont know? Maybe their helping someone else?

OpenStudy (blues):

Bugger. In that case, I shall wait until they return to finish the prob.

OpenStudy (anonymous):

Okay, I'm going to go have a look to see if I got any more help on my question. @blues you any good with essays by any chance? Because I desperately need help with my question. @jatt3092, If you need any more help, just leave me a message or tag me in a comment. :)

OpenStudy (anonymous):

yeah sorry guys I was busy another problem my apologizes..but yeah shay whats ur essay question i could help u out.

OpenStudy (blues):

Hey Jatt, do you know how to get from the renormalized proportions of homozygous dominant individuals and heterozygous individuals to new values for p and q?

OpenStudy (anonymous):

@jatt3092, its is posted right here: http://openstudy.com/users/shay88#/updates/4f81f0e8e4b0505bf08380d4 I will let you figure out @blue's question on your own. Just think about all the stuff that you've been taught and the formulas.

OpenStudy (anonymous):

dont' we have to figure out the proportion of homozygous recessive inviduals?

OpenStudy (anonymous):

so basically once we find the new question you just have to square it and u got the answer

OpenStudy (anonymous):

q value* not question

OpenStudy (blues):

Yes, exactly. You need to find new values for p and q for the post die off population. Once you have found q, you have to square it to get the number of homozygous recessive individuals in the new population. But before that you have to solve for p and q...

OpenStudy (anonymous):

Well that's for before the catalysmic. You have to find the new q value using the p^2 value we found. Do you have any ideas of how to go about this?

OpenStudy (anonymous):

no :(

OpenStudy (blues):

OK Jatt, when we renormalized the population of surviving heterozygotes and homozygous dominant individuals, we found the proportion of pp and qp individuals in the new population. That is what the new pp = 0.5384 and new 2pq = 0.46154 are all about. Is that clear?

OpenStudy (anonymous):

Use another formula given to you for the ALLELE frequencies, instead of the one given for the population frequencies that we just used.

OpenStudy (blues):

Shay, I'm sorry but I don't know which way you're going with that last.

OpenStudy (anonymous):

Well we have p^2 so you square that to get p. Right? You following so far?

OpenStudy (anonymous):

yes im clear about those

OpenStudy (anonymous):

Then we need q for the recessive allele frequency.

OpenStudy (anonymous):

blue is the answer 0.0729

OpenStudy (anonymous):

The equation for population frequencies is p^2 + 2pq = q^2 The equation for allele frequencies though is different.

OpenStudy (anonymous):

Its not by my calculations jatt

OpenStudy (blues):

So essentially those are two new equations for p and q. We have: p^2 = 0.5384 2pq = 0.4615. We can solve those very easily for the new allele frequencies p and q, can't we? p = sqrt(0.5384) = 0.7338 Once we have that we can get q from the second equation: q = 0.4615 / (2*p) q = 0.4615 / (2*0.7338) q = 0.3145. And then to get the frequency of homozygous recessive (qq) individuals in the new population is trivial. We just compute q^2. q ^ 2 = 0.3145 ^ 2 = 0.0989. Clear?

OpenStudy (anonymous):

Well there is an easier way to do it using the allele frequency equation which is p + q = 1. So once you've got p = 0.7338, you get q by manipulating that, and getting q = 1- p. And then squaring that value.

OpenStudy (blues):

Or you can do it that way if you please. ;D

OpenStudy (anonymous):

So I get 0.0729. My mistake, you were right @jatt3092

OpenStudy (anonymous):

blue u got a different answer :S

OpenStudy (anonymous):

Or no 0.0709. Sorry Hold on I'm just going to go grab an actual calculator instead of the computer one and check this out.

OpenStudy (blues):

I double checked my math and I actually have to get going. Post it and I'll catch up with it. :D

OpenStudy (anonymous):

k thanks you :)

OpenStudy (anonymous):

I got 0.2662 for my q value, so that might be the problem.

OpenStudy (anonymous):

nah blues did it right

OpenStudy (anonymous):

Yeah I cant figure out what I did wrong, but when you add blues values together for p^2+ 2pq + q^2 it comes up for 1. So ignore me a little bit. and go with what blues did.